OpenStudy (anonymous):

Javier designs leather shoes in his workshop. He charges $45 for women's shoes and $50 for mens shoes. He needs to make $2,500 per week to meet his expenses. This situation can be given by 45 X plus 50 y equals 2500, where X is the number of women shoes and y is the number of mens shoes designed. If he does not have any female customers during a particular week, how many male customers does he need to meet his expenses? (X > or equal to 0, y > or equal to 0, and y take only integer values)

OpenStudy (mathstudent55):

If he has no female customers, then x = 0. 45x + 50y = 2500 If you let x = 0, this becomes: 45(0) + 50y = 2500 50y = 2500 Solve for y to find out how many men's pairs of shoes he needs to sell.

OpenStudy (anonymous):

45 (0) + 50 (50) = 2500??

OpenStudy (mathstudent55):

Correct, so the answer is 50 men's shoes.